An atom has only has two absorption lines, at 300 nm and 450 nm. What is the wavelength of the one line in the emission spectrum which does not appear in the absorption spectrum.
Answer #1

We have the wavelength which does not appear in the absorption spectrum. as, \frac{1}{\lambda} = \frac{1}{\lambda_1} - \frac{1}{\lambda_2}

\Rightarrow \frac{1}{\lambda} = \left [\frac{1}{300} - \frac{1}{450} \right ]\frac{1}{10^{-9}}

\Rightarrow \frac{1}{\lambda}= 1.11 10^{-3} * * 10^{9} = 1.11 * 10^{6}

\Rightarrow \lambda =\frac{1}{ 1.11 * 10^{6}} = 900 nm